OpenStudy (anonymous):

The slope of the line joining the coordinate points (3, -1) and (-4, 7) is -

OpenStudy (jhannybean):

In finding the slope, use the formula \[\large \frac{ \color{red}{y _{2}-y _{1}} }{ \color{green}{x _{2}-x _{1}} }\]

OpenStudy (anonymous):

slope of the line joining two points (x1,y1) and (x2,y2) is given by (y2-y1)/(x2-x1)

OpenStudy (anonymous):

here x1=3 and y1=-1 and x2=-4 and y2=7

OpenStudy (jhannybean):

So you would have \[\large\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }= \frac{ 7-(-1) }{ -4-3 }\] now you can solve.